# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ 3.1 Randomized Matrix Markov Inequality
where $\mathbf{X}$ is an $\mathcal{S}_{d}^{+}$-valued random matrix and $\mathbf{A}\in\mathcal{S}_{d}^{++}$ is deterministic.
To achieve this, let us first define what it means for a random matrix to be uniformly and super-uniformly distributed. Note that it is easy to see that the randomizer $U$ in (UMI) and (UVI) can be replaced by any random variable that is stochastically larger; that is, one satisfying the following definition:
Definition 3.1.
We call a positive random scalar $U$ super-uniform if $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$; that is, $U$ is stochastically greater than or equal to the uniform distribution $\operatorname{Unif}_{(0,1)}$.
This has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
Definition 3.2.
Straightforwardly the identity matrix $\mathbf{I}$ is spectrally (hence trace) super-uniform. If $U\sim\operatorname{Unif}_{(0,1)}$, $U\mathbf{I}$ is $\mathcal{S}_{d}^{[0,1]}$-spectrally uniform, and $\left\{\mathbf{v}\mathbf{v}^{\mathsf{T}}:\|\mathbf{v}\|\leqslant 1\right\}$-trace uniform. Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (UMI).
Theorem 3.3 (Uniformly Randomized Matrix Markov Inequality).
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (UMMI)
Proof.
By our assumptions,
[…]
concluding the proof. ∎
[…]
The following is an example where the inequality of (UMMI) holds with equality. We take $\mathbf{U}=U\mathbf{I}$ where $U\sim\operatorname{Unif}_{(0,1)}$, and $\mathbf{A}\in\mathcal{S}_{d}^{++}$. Let $\mathbf{x}$ be a random vector in the ellipsoid $\{\mathbf{x}:\mathbf{x}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x}\leqslant 1\}\subseteq\mathbb{R}^{d}$. Then,
where the second equality above is (UMMI) applied on the random matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$. It is not hard to see here the condition for equality is met. Intuitively, this hints that (UMMI) is tighter when the random matrix $\mathbf{X}$ is approximately of lower rank, a point that we shall return to later in this paper with statistical examples.
3.2 Randomized Matrix Chebyshev Inequalities
Recall that the variance of a random symmetric matrix $\mathbf{X}$ is defined simply as $\mathbf{Var}(\mathbf{X})=\mathbb{E}(\mathbf{X}-\mathbb{E}\mathbf{X})^{2}$. Applying Theorem 3.3 to the square of one square-integrable matrix and the square of the sum of several i.i.d. square-integrable matrices, we have the following randomization of the one-matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14],
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$ (12)
@@ 3.2 Randomized Matrix Chebyshev Inequalities
and the implied law of large numbers for the average $\overline{\mathbf{X}}_{n}$ of $n$ i.i.d. symmetric matrices with mean $\mathbf{M}$ and variance $\mathbf{V}$ by Ahlswede and Winter [2002, Theorem 15],
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (MCI)
Lemma 3.4 (Uniformly Randomized One-Matrix Chebyshev Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2}).$ (13)
Proof.
By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})$. By Theorem 3.3, we have $\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$. ∎
We can immediately apply Lemma 3.4 to multiple matrices, obtaining the following weak-law-of-large-numbers style inequality that shrinks when $n$ gets larger. Recall that the anticommutator between two sqaure matrices $\mathbf{A},\mathbf{B}$ is defined as $\{\mathbf{A},\mathbf{B}\}=\mathbf{A}\mathbf{B}+\mathbf{B}\mathbf{A}$.
Theorem 3.5 (Uniformly Randomized $n$-Matrix Chebyshev Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all. Then, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, we have
$\mathbf{Var}(\overline{\mathbf{X}}_{n})\preceq n^{-1}\mathbf{V}.$ (14)
Consequently, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (UMCI)
Proof.
Note that
$\small\mathbf{Var}(\overline{\mathbf{X}}_{n})=n^{-2}\mathbb{E}\left(\sum_{i=1}^{n}(\mathbf{X}_{i}-\mathbf{M})\right)^{2}=n^{-2}\left(\sum_{i=1}^{n}\mathbf{Var}(\mathbf{X}_{i})+\sum_{i<j}\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}\right)\preceq n^{-1}\mathbf{V},$ (15)
and apply Lemma 3.4 to conclude the proof. ∎
It is easy to see that (UMCI) recovers as special cases (MCI) and (CI).
Remark 3.6.
The assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
For $p\in[1,2)$, a $p$th moment bound can lead to an infinite variance version of (13), whose non-randomized ($\mathbf{U}=\mathbf{I}$) version is mentioned by Tu [2017].
Lemma 3.7 (Uniformly Randomized One-Matrix $p$-Chebyshev Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$ that has a finite $p$th central moment $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}$ and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (16)
Proof.
By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})$. By Theorem 3.3, we have $\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p})$. ∎
However, an $L^{p}$ analogy of (UMCI) (i.e. a matrix version of the $p$-Chebyshev inequality (5)) is not straightforward. This is largely because while the contraction of variance rule (14) exists for matrices, we do not know of any matrix version of $p$th moment contraction rule. In fact, the scalar $p$th moment contraction rule was only recently established (by Cherapanamjeri et al. [2022, Lemma 4.2] and Wang et al. [2021, Lemma 7], see the proof of (5) in Section A.1). Section A.2 contains a conjectured result along these lines, as well as a looser result that is the best we can currently achieve.
3.3 Randomized Matrix Chernoff Bounds
We now establish the classical Chernoff-type bounds with randomization for light-tailed random matrices that have finite moment generating functions (MGFs).
Lemma 3.8 (Uniformly Randomized One-Matrix Chernoff Bound).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}$ and real number $\gamma$,
$\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}}).$ (17)
Proof.
By Fact 2, we have $\mathbb{P}\left(2\gamma\mathbf{X}\npreceq\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$. By Theorem 3.3, we have $\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}})$, concluding the proof. ∎
The seemingly indigestible tail threshold $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$ in (17) can be largely simplified in any of the following scenarios. First, one can simply take the trace on the tail event above, obtaining due to Fact 1 and the monotonicity of the trace,
$\mathbb{P}\left(\mathrm{tr}\mathbf{X}>\mathrm{tr}\mathbf{A}+\frac{1}{2\gamma}\mathrm{tr}\log\mathbf{U}\right)\leqslant\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right),$ (18)
without any additional assumptions; Second, if $\mathbf{U}=U\mathbf{I}$ where $U$ is scalar super-uniform, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=\mathbf{A}+\frac{1}{2\gamma}\mathbf{I}\log U$; Third, if $\mathbf{A}=a\mathbf{I}$, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=a\mathbf{I}+\frac{1}{2\gamma}\log\mathbf{U}$.
It is known that, unlike the scalar case, the one-matrix MGF is not linearly additive. Once again, the complication is partially resolved when the trace is applied.
Fact 5 (Lemma 3.4 of Tropp [2012]).
Let $\mathbf{Z}_{1},\dots,\mathbf{Z}_{n}$ be independent random matrices taking values in $\mathcal{S}_{d}$. Then,
$\mathrm{tr}\left\{\mathbb{E}\exp\left(\sum\mathbf{Z}_{i}\right)\right\}\leqslant\mathrm{tr}\exp\left(\sum\log\mathbb{E}\mathrm{e}^{\mathbf{Z}_{i}}\right).$ (19)
This implies the following bound that recovers as special cases both the scalar Chernoff-Hoeffding inequality (6) when taking $d=\mathbf{U}=1$, and the “master tail bound” on $\lambda_{\max}(\overline{\mathbf{X}}_{n}-\mathbf{M})$ by Tropp [2012, Theorem 3.6] when taking $\mathbf{U}=\mathbf{I}$.
Theorem 3.9 (Uniformly Randomized Matrix Chernoff-Hoeffding Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and matrix-valued MGF $\mathbf{G}(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all. Let $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. For any $a,t>0$,
$\mathbb{P}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\npreceq a\mathbf{I}+\frac{\log\mathbf{U}}{\gamma}\right)\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\mathrm{e}^{-\gamma a}.$ (20)
Proof.
By Fact 5, $\mathrm{tr}(\mathbb{E}\mathrm{e}^{\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})})\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$. Using Lemma 3.8 on $\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})$ with $\mathbf{A}=a\mathbf{I}$ concludes the proof. ∎
Again, the bound remains valid if only an upper bound of the matrix MGF $\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\mathbf{G}(\gamma)$ or (more commonly) the weaker log-MGF bound $\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\log\mathbf{G}(\gamma)$ is known. Table 1 contains some important known matrix (log-)MGFs (see Tropp [2012, 2015]). An asymmetric Hoeffding bound is missing here but shall be discussed in Section 5 later since it requires some further techniques beyond the matrix log-MGF.
Name Condition $\mathbf{G}(\gamma)$ $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$
Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $=\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
Bennett I $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})=\mathbf{V}$ $\preceq\mathrm{e}^{(\mathrm{e}^{\gamma}-\gamma-1)\mathbf{V}}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
Bennett II $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})\preceq\mathbf{V}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
Symmetric Hoeffding $(\mathbf{X}_{1}-\mathbf{M})^{2}\preceq\mathbf{B},\ \mathbf{X}_{1}-\mathbf{M}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}-\mathbf{X}_{1}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{B}/2n})$
Table 1: Some important matrix MGFs and log-MGFs. Here $\mathbf{B},\mathbf{C},\mathbf{V}$ are deterministic matrices, and =$\stackrel{{\scriptstyle d}}{{=}}$ means two random matrices having the same distribution. The proofs are in Section B.1.
4 Maximal Inequalities of Matrix (Super-/Sub-)Martingales
Some interesting tail bounds are obtained if we consider matrix-valued supermartingales and submartingales and their maximal inequalities. These processes can be defined on forward or backward filtrations and in both these cases, well-known scalar-valued maximal inequalities can be extended to matrices.
4.1 Matrix-Valued Supermartingales
@@
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued stochastic process adapted to the filtration $\{\mathcal{F}_{n}\}$. [^3] We shall often omit the filtration $\{\mathcal{F}_{n}\}$ when unambiguous from the context. We say that $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$. Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$). See also a “dominated integrability” lemma we prove in Section B.2.
@@
An $\mathcal{S}_{d}$-valued process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every deterministic vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
Proof.
$\{\mathbf{Y}_{n}\}$ is a supermartingale (21)
$\iff$ For all $\mathbf{v}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (25)
This completes the proof. ∎
This leads to an analog of Doob’s martingale convergence theorem for positive semidefinite matrix supermartingales.
@@
For any $\mathbf{v}\in\mathbb{R}^{d}$, the process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a scalar-valued nonnegative supermartingale, which converges a.s. to an integrable random scalar due to Doob’s martingale convergence theorem (see e.g. Corollary 11.5 in Klenke [2013]). First, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s diagonal converges to integrable scalars. Then, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that the rest of $\mathbf{Y}_{n}$’s entries also converge to integrable scalars. The limit $\mathbf{Y}_{\infty}=\lim_{n\to\infty}\mathbf{Y}_{n}$ is in $\mathcal{S}_{d}^{+}$ because $\mathcal{S}_{d}^{+}$ is a closed set. Finally, for any $\mathbf{v}\in\mathbb{R}^{d}$,
@@
Next, we see that using a similar trick, any scalar version of the optional stopping (sampling) theorems extends to symmetric matrices as well. For example:
Lemma 4.3 (Matrix Optional Stopping).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then,
@@
    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale on $\{\mathcal{F}_{n}\}$.

### v4

@@ 2.3.2 Ahlswede-Winter Bounds
where $\mathbf{X}$ is an $\mathcal{S}_{d}^{+}$-valued random matrix and $\mathbf{A}\in\mathcal{S}_{d}^{++}$ is deterministic. This leads to the direct corollary of a matrix Chebyshev inequality that serves as a weak law of large numbers for the average $\overline{\mathbf{X}}_{n}$ of $n$ i.i.d. symmetric matrices with mean $\mathbf{M}$ and (matrix) variance $\mathbf{V}=\mathbb{E}(\mathbf{X}-\mathbf{M})^{2}$ [Ahlswede and Winter, 2002, Corollary 16],
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (10)
Ahlswede and Winter [2002] also attempt to use this “un-scalarized” method to generalize the Cramér-Chernoff method for scalars. However, as summarized by [Tropp, 2012, Section 3.7], this leads to suboptimal bounds that involve “the sum of eigenvalues” as opposed to “the eigenvalue of a sum” which the sharper bounds (obtained by “scalarizing” with Lieb’s concavity theorem as mentioned in Section 2.3.1) involve.
Our time-uniform, martingale-based main results are often obtained via (MMI). In particular, we show that this Loewner-ordered approach pioneered by Ahlswede and Winter [2002] does lead to sharp matrix concentration inequalities and tests, fixed-time and time-uniform alike, comparable to and sometimes better than the results cited in Section 2.3.1. We provide a self-contained proof of (MMI) in Appendix A that also comes with the extension of randomization.
3 Forward PSD Supermartingales
We consider in this section a probability space with a usual forward filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$. That is, $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\dots\subseteq\mathcal{F}_{\infty}$. Within the scope of this section, we shall often omit the filtration $\{\mathcal{F}_{n}\}$ when unambiguous from the context, calling a scalar- or matrix-valued process “adapted” if the $n$-indexed element is $\mathcal{F}_{n}$-measurable; “predictable” if $\mathcal{F}_{n-1}$-measurable. Stopping times are all with respect to $\{\mathcal{F}_{n}\}$.
3.1 Convergence, Optional Stopping, and Maximal Inequalities
@@
Recall our definition of a matrix-valued supermartingale from the beginning of the paper that an adapted $\mathcal{S}_{d}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$. Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$). See also a “dominated integrability” lemma we prove in Section B.1.
@@
An $\mathcal{S}_{d}$-valued adapted process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every non-random vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
The very straightforward proof of Lemma 3.1 can be found in Section B.2. It leads to an analog of Doob’s martingale convergence theorem for PSD matrix supermartingales.
@@
For any $\mathbf{v}\in\mathbb{R}^{d}$, the process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a scalar-valued nonnegative supermartingale, which converges a.s. to an integrable random scalar due to Doob’s martingale convergence theorem (see e.g. Corollary 11.5 in Klenke [2013]). First, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s diagonal converges to integrable scalars. Then, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s off-diagonal entries also converge to integrable scalars. The limit $\mathbf{Y}_{\infty}=\lim_{n\to\infty}\mathbf{Y}_{n}$ is in $\mathcal{S}_{d}^{+}$ because $\mathcal{S}_{d}^{+}$ is a closed set. Finally, for any $\mathbf{v}\in\mathbb{R}^{d}$,
@@
The convergence of PSD supermartingales ensures that we can speak of the stopped matrix $\mathbf{Y}_{\tau}$ for any stopping time $\tau$, even when $\mathbb{P}(\tau=\infty)>0$. We can use a similar method to extend some scalar versions of the optional stopping (sampling) theorems to matrices. For example:
Theorem 3.3 (Matrix Optional Stopping).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time. Then,
@@
    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale.